onsider the following information about three stocks:

State of
Economy
Probability of State of Economy Rate of Return if State Occurs
Stock A Stock B Stock C
Boom 0.20 0.20 0.32 0.54
Normal 0.45 0.18 0.16 0.14
Bust 0.35 0.02 −0.34 −0.42

a-1. If your portfolio is invested 40% each in A and B and 20% in C, what is the portfolio expected return? (Do not round intermediate calculations. Enter the answer as a percent rounded to 2 decimal places.)

Portfolio expected return             %

a-2. What is the variance? (Do not round intermediate calculations. Round the final answer to 8 decimal places.)

Variance            

a-3. What is the standard deviation? (Do not round intermediate calculations. Enter the answer as a percent rounded to 2 decimal places.)

Standard deviation             %

b. If the expected T-bill rate is 4.00%, what is the expected risk premium on the portfolio? (Do not round intermediate calculations. Enter the answer as a percent rounded to 2 decimal places.)

Expected risk premium             %

c-1. If the expected inflation rate is 2.00%, what are the approximate and exact expected real returns on the portfolio? (Do not round intermediate calculations. Enter the answers as a percent rounded to 2 decimal places.)

Approximate expected real return %
Exact expected real return %

c-2. What are the approximate and exact expected real risk premiums on the portfolio? (Do not round intermediate calculations. Enter the answers as a percent rounded to 2 decimal places.)

Approximate expected real risk premium %
Exact expected real risk premium %
Answer #1

A B с D E F G H J K 1 2 3 4 5 6 7 8 Expected Returns 9 Variance 10 Standard Deviation 11 Expected Risk Premium 12 Approximate

A B C D E F H 1 1 ProbA 0.2 0.2 0.45 0.18 (R-Mean)^2 B = (C5-C$8)^2 0.32 =(C6-C$8)^2 0.16 = (C7-C$8)^2 -0.34 =SUMPRODUCT($B$5